Blue Elk Manufacturing has the following end-of-year balance sheet:

Blue Elk Manufacturing

Balance Sheet

For the Year Ended on December 31

Assets Liabilities
Current Assets: Current Liabilities:
Cash and equivalents $150,000 Accounts payable $250,000
Accounts receivable 400,000 Accrued liabilities 150,000
Inventories 350,000 Notes payable 100,000
Total Current Assets $900,000 Total Current Liabilities $500,000
Net Fixed Assets: Long-Term Bonds 1,000,000
Net plant and equipment(cost minus depreciation) $2,100,000 Total Debt $1,500,000
Common Equity
Common stock 800,000
Retained earnings 700,000
Total Common Equity $1,500,000
Total Assets $3,000,000 Total Liabilities and Equity $3,000,000

The firm is currently in the process of forecasting sales, asset requirements, and required funding for the coming year. In the year that just ended, Blue Elk Manufacturing generated $350,000 net income on sales of $14,000,000. The firm expects sales to increase by 18% this coming year and also expects to maintain its long-run dividend payout ratio of 45%.

Suppose Blue Elk Manufacturing’s assets are fully utilized. Use the additional funds needed (AFN) equation to determine the increase in total assets that is necessary to support Blue Elk Manufacturing’s expected sales.

$567,000

$486,000

$540,000

$459,000

When a firm grows, some liabilities grow spontaneously along with sales. Spontaneous liabilities are a source of capital that the firm will generate internally, so they reduce the need for external capital. How much of the total increase in assets will be supplied by spontaneous liabilities for Blue Elk Manufacturing this year?

$61,200

$64,800

$72,000

$75,600

In addition, Blue Elk Manufacturing is expected to generate net income this year. The firm will pay out some of its earnings as dividends but will retain the rest for future asset investment. Again, the more a firm generates internally from its operations, the less it will have to raise externally from the capital markets. Assume that the firm’s profit margin and dividend payout ratio are expected to remain constant.

Given the preceding information, Blue Elk Manufacturing is expected to generate from operations that will be added to retained earnings.

According to the AFN equation and projections for Blue Elk Manufacturing, the firm’s AFN is.
Answer #1
EFN is given by the equation:
EFN = (A/S) x (Δ Sales) - (L/S) x (Δ Sales) - (PM x FS x (1-d))
where,
A / S = Assets that change givendirectly with sales.
Δsales = Change in sales
L / S = Liabilities that change directly with sales
PM = Profit Margin on Sales = net income / sales.
FS = Forecasted Sales
d = dividend payout ratio
(1 - d) = retention ratio
1] Increase in total assets = (3000000/14000000)*14000000*18% = $           5,40,000
2] Increase in spontaneous liabilities = (400000/14000000)*14000000*18% = $              72,000
3] Retention = 14000000*118%*2.5%*(1-45%) = $           2,27,150
[PM ratio = 350000/14000000 = 2.50%]
4] EFN = 540000-72000-227150 = $           2,40,850